### v2

# A unified weighting framework for evaluating nearest neighbour classification

Abstract
We present the first comprehensive and large-scale evaluation of classical (NN), fuzzy (FNN) and fuzzy rough (FRNN) nearest neighbour classification. We standardise existing proposals for nearest neighbour weighting with kernel functions, applied to the distance values and/or ranks of the nearest neighbours of a test instance. In particular, we show that the theoretically optimal Samworth weights converge to a kernel. Kernel functions are closely related to fuzzy negation operators, and we propose a new kernel based on Yager negation. We also consider various distance and scaling measures, which we show can be related to each other. Through a systematic series of experiments on 85 real-life classification datasets, we find that NN, FNN and FRNN all perform best with Boscovich distance, and that NN and FRNN perform best with a combination of Samworth rank- and distance-weights and scaling by the mean absolute deviation around the median ($r_{1}$), the standard deviation ($r_{2}$) or the semi-interquartile range ($r_{\infty}^{*}$), while FNN performs best with only Samworth distance-weights and $r_{1}$- or $r_{2}$-scaling. However, NN achieves comparable performance with Yager-$\frac{1}{2}$ distance-weights, which are simpler to implement than a combination of Samworth distance- and rank-weights. Finally, FRNN generally outperforms NN, which in turn performs systematically better than FNN.
keywords:
classification , fuzzy nearest neighbours , fuzzy negation , fuzzy rough nearest neighbours , kernels , nearest neighbours , weighting
[…]
1 Introduction
Nearest neighbours (NN) [1] is one of the oldest and most widely used classification algorithms. Despite the development of other successful classifiers, it remains an attractive option in many settings due to its simplicity [2] and good performance for many problems, especially when decision boundaries are irregular [3]. NN and its variants are applied in such diverse domains as healthcare [4, 5, 6], agriculture [7, 8], geology [9, 10], industry [11, 12, 13] and natural language processing [14, 15, 16], and have also been adapted to big data settings [17, 18, 19].
While NN is a relatively simple algorithm, it still requires setting a few hyperparameters. Even in its most basic form, we need to choose a distance measure and a method to rescale the attributes of the dataset. In addition, it is generally advisable to choose the number of neighbours $k\geq 1$ on which predictions are based. Moreover, there have been many proposals in the literature to weigh the contribution of the $k$ nearest neighbours of a test record differently.
In principle, it is possible to resolve these choices for any given problem by picking the combination of hyperparameter values that performs best on cross-validated training data. However, in practice, it is often more convenient to focus these efforts on the value $k$, and set the other hyperparameters to values that are known to be good enough. To aid this approach, it would be useful to have an idea which choices generally perform better than others.
Building on the traditional form of (weighted) NN described above, some authors have proposed further-reaching modifications that incorporate fuzzy set theory: fuzzy nearest neighbours (FNN) [20] and fuzzy rough nearest neighbours (FRNN) [21]. FNN operates in a similar way to NN, but uses fuzzified class membership degrees of training records, while FRNN models each decision class as a fuzzy set, and calculates the membership degrees of a test record in these fuzzy sets. Like NN, these algorithms require a choice of distance measure, scale, number of neighbours $k$ and weighting scheme.
While there has been no shortage of new ideas for NN variants, it remains largely unclear which proposals work better in practice. The goal of this paper is to address this. Concretely, our contributions are the following:
  • 1.
    We present a comprehensive overview of the different weighting proposals for NN, FNN and FRNN in the literature.
  • 2.
    We establish a common weighting framework in terms of kernel functions that unifies these proposals.
  • 3.
    In particular, we prove that the theoretically optimal rank-weights identified by Samworth [22] converge to a specific kernel function as $k$ increases.
  • 4.
    We show that kernel functions are closely related to fuzzy negation operators. In particular, the weights proposed by Gou et al. [23] correspond to Sugeno negation. Inspired by this, we propose our own weighting kernel based on Yager negation, which has a distinct contour from existing weighting proposals.
  • 5.
    We introduce the concept of the Minkowski $p$-radius $r_{p}$ of a dataset, and show that $r_{2}$ and $r_{\infty}$ are, respectively, the standard deviation and the half-range, two commonly used measures of dispersion. We propose that $r_{1}$, the mean absolute deviation around the median, should also be considered as a scaling measure, in particular because it is less sensitive to outliers.
  • 6.
    We conduct a large-scale experiment on 85 real-life classification datasets, comparing distance and scaling measures and weighting kernels. To the best of our knowledge, this is the first large-scale evaluation of all NN weighting proposals, and the first direct comparison of FRNN with NN and FNN.
In Section 2, we present an overview of the literature on nearest neighbour classification variants. Next, we present our own proposals in Section 3. We then describe our experimental setup (Section 4) and present the results (Section 5), before concluding (Section 6).
2 Background
In this section, we briefly review the existing literature on nearest neighbour weighting, as well as previous experimental comparisons.
2.1 Weighted nearest neighbour classification
Figure 1: Illustration of NN classification with $k=3$. The class scores of a test record are calculated on the basis of its $k$ nearest training records.
In the early literature, nearest neighbour prediction (Fig. 1) arose as a form of non-parametric (or distribution-free) statistical estimation, and was generally referred to as such. It was first formally presented in 1951 for classification, by Fix & Hodges [1]. The idea to weigh the contribution of neighbours differently was initially proposed for regression, perhaps first by Watson [24], Royall [25] and Shepard [26]. This was inspired by an earlier idea to estimate the value of a density function in a point as a weighted sum, with weights corresponding inversely to the distances to the sample observations [27]. Dudani [28, 29] appears to have been the first to propose weighted nearest neighbours for classification.
We can formally define weighted nearest neighbour classification as follows. Let $d$ be a distance measure and $k$ a positive integer, then the score for a decision class $C$ and a test record $y$ is:
$\left.\sum_{i\leq k|x_{i}\in C}s_{i}\middle/\sum_{i\leq k}s_{i}\right.,$ (1)
where $x_{i}$ is the $i$th nearest neighbour of $y$ in the training set $X$ (according to $d$), and $s_{i}$ the weighted vote assigned to $x_{i}$, which remains to be defined. In practice, all proposals define $s_{i}$ in terms of the distance $d_{i}$ between $y$ and $x_{i}$ and/or the rank $i$. We will refer to these strategies as, respectively, distance- and rank-weighting. We recover classical unweighted nearest neighbour classification by choosing constant $s_{i}$, e.g. $s_{i}=1$.
For rank-weighting, there have been proposals that let the weights depend linearly on the rank [25, 28, 29, 30], quadratically [30, 31], reciprocally linearly [32], and according to the Fibonacci sequence [33]. Relatively recently, Samworth [22] has established theoretically optimal weights:
$s_{i}=\frac{1}{k}\left(1+\frac{m}{2}-\frac{m}{2k^{\frac{2}{m}}}\left(i^{1+\frac{2}{m}}-(i-1)^{1+\frac{2}{m}}\right)\right),$ (2)
where $m$ is the dimensionality of the attribute space.
Proposals for distance-weighting have included weights that depend linearly on distance [24, 28, 29], reciprocally linearly [28, 29] and reciprocally quadratically [26]. Inspired by the work of Shepard [34], Zavrel [35] has proposed Laplacian weights of the form $e^{-d_{i}}$.
The linear distance-weights given by Dudani [28, 29] take the following form:
$s_{i}=\begin{dcases}\frac{d_{k}-d_{i}}{d_{k}-d_{1}}&k>1;\\ 1&k=1.\\ \end{dcases}$ (3)
For these weights, Dudani demonstrated a lower classification error than unweighted NN on a synthetic dataset. However, Bailey and Jain [36] subsequently showed that this was due to the fact that Dudani had counted all ties as errors, and that when these are resolved instead (e.g. by randomly choosing a class), the performance of weighted and unweighted NN was similar on the synthetic dataset. Moreover, Bailey and Jain also proved that the asymptotic classification error of unweighted NN is minimal among all possible weighted variants of NN. This in turn elicited a response by Macleod et al. [37], who argued that there exist finite classification problems where some distance-weighted variants of NN do have lower error. In order to demonstrate this, they used the following modified weights, which address the fact that in (3), the $k$th weight is always 0 (if $k>1$):
$s_{i}=\begin{dcases}\frac{d_{k}-d_{i}+d_{k}-d_{1}}{2(d_{k}-d_{1})}&k>1;\\ 1&k=1.\\ \end{dcases}$
Another modification of Dudani’s linear weights was proposed by Gou et al. [23]:
$s_{i}=\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{d_{k}+d_{1}}{d_{k}+d_{i}}$
Finally, Gou et al. [32] have proposed a weighting scheme that combines the linear distance-weights of Dudani with reciprocal rank-weights:
$s_{i}=\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{1}{i}$
2.2 Kernel weighting
Nearest neighbour prediction is closely related to another form of non-parametric estimation, in which the prediction for a test record is a weighted sum of training values (e.g. [38]). The weights are determined by a so-called kernel function that is applied to the distances between the test record and the training records. By choosing a decreasing kernel function, nearby training records receive greater weight, and by choosing a kernel function with finite support, the prediction is effectively limited to neighbours within a fixed distance, although the number of neighbours will generally vary for different test records.
There have been several proposals to extend the use of kernel functions to nearest neighbour-weighting. A very early proposal by Royall [25] applies a linear (or triangular) kernel to the scaled rank $\frac{i}{k}$, while Altman [31] suggests the use of a quadratic kernel. Both Wilson & Martinez [39, 40] and Hechenbichler & Schliep [41] apply kernel functions to nearest neighbour distances, rescaled to values in $[0,1]$ by dividing by the $k$th distance. In addition to constant and linear kernels, Wilson & Martinez use Laplacian and Gaussian kernels, while Hechenbichler & Schliep consider a biquadratic kernel $a\mapsto\frac{15}{16}(1-a^{2})^{2}$.
2.3 Fuzzy nearest neighbour classification
Figure 2: Illustration of FNN classification with $k=3$. The class scores of a test record are calculated on the basis of its $k$ nearest training records, whose class memberships are first fuzzified on the basis of their own $k$ nearest training records.
There have been many proposals to modify nearest neighbour classification with fuzzy set theory [42]. The most prominent of these is the fuzzy nearest neighbours (FNN) classifier (Fig. 2) of Keller et al. [20]. It defines the membership of a test record $y$ in the decision class $C$ as
$\left.\sum\limits_{i\leq k}u_{i}\cdot 1/d_{i}^{2/(q-1)}\middle/\sum\limits_{i\leq k}1/d_{i}^{2/(q-1)}\right.,$ (4)
for a choice of $q>1$, where $d_{i}$ is the distance between $y$ and its $i$th nearest neighbour $x_{i}$, and $u_{i}$ is the class membership of $x_{i}$ in $C$. Keller et al. proposed two different options for $u_{i}$. Either $u_{i}$ can be chosen to be the crisp class membership of $x_{i}$ in $C$, or it can be fuzzified as follows:
$u_{i}=\begin{dcases}0.51+0.49\cdot n_{C}(x_{i})/k&\text{if }x_{i}\in C;\\ 0.49\cdot n_{C}(x_{i})/k&\text{if }x_{i}\notin C;\\ \end{dcases},$ (5)
where $n_{C}(x_{i})$ is the number of neighbours of $x_{i}$ that belong to $C$, from among its $k$ nearest neighbours.
The original motivation for FNN given by Keller et al. was twofold. Firstly, by providing class scores (4) rather than a crisp prediction that the test record $y$ belongs to one class or another, FNN expresses the ‘strength’ of its class membership. Secondly, FNN lets the contribution of the nearest neighbours of $y$ depend both on their distance to $y$ and, when using the fuzzified class membership (5), on how ‘typical’ they are for a given decision class. However, the ability to provide class scores does not actually set it apart from classical NN (1). Likewise, we saw in Subsection 2 that there is a long tradition of letting the predictions of NN depend on the distance of the neighbours of $y$. In Subsection 3.4, we will further analyse how FNN is related to NN.
2.4 Fuzzy rough nearest neighbour classification
Figure 3: Illustration of FRNN classification with $k=3$. The class scores of a test record are calculated on the basis of its $k$ nearest training records within each decision class (upper approximation) as well as its $k$ nearest training records in the complement of each decision class (lower approximation).
A more fundamentally different proposal has come in the form of fuzzy rough nearest neighbour (FRNN) classification (Fig. 3), originally proposed by Jensen & Cornelis [21]. This is based on fuzzy rough sets [43], a fuzzified variant of rough sets [44]. For each decision class $C$, we define two fuzzy sets, its upper approximation $\overline{C}$ and its lower approximation $\underline{C}$, as well as their mean, and the membership of a test record $y$ in any one of these can be used as a class score. A weighted variant of fuzzy rough sets was first introduced by Cornelis et al. [45], and we use here the updated formulation of FRNN presented in [18].
Let $d$ be a distance measure and $k$ a positive integer, then the membership of a test record $y$ in $\overline{C}$ and $\underline{C}$ is defined, as, respectively:
[…]
where $d^{+}_{i}$ and $d^{-}_{i}$ are the $i$th nearest neighbour distance of $y$ in, respectively, $C$ and $X\setminus C$, and $w_{i}$ is a weight which depends on the rank $i$ and which remains to be defined.
Previous proposals for $w$ have included weight vectors that are constant [45] or that depend linearly [46], reciprocally [47] or exponentially [45] on the rank $i$ (see [48] for an overview).
FRNN was originally proposed by Jensen & Cornelis [21] in reaction to a previous proposal by Sarkar [49] to modify FNN with a so-called fuzzy rough ownership function. In contrast to the proposal by Sarkar, FRNN is not a modification of FNN, but uses the core concepts of the upper and lower approximation from fuzzy rough set theory to obtain classification scores. However, a common motivation for both proposals is the observation (originally by Sarkar) that the class membership values produced by FNN automatically sum to one, and that it therefore cannot identify situations where we do not have enough information to make a reliable prediction for the test record $y$, i.e. when $y$ doesn’t particularly resemble any of the training records. In such a situation, FRNN predicts low upper approximation values and high lower approximation values for all decision classes.
2.5 Previous experiments
Despite the extensive literature on NN classification, there have only been a small number of experimental evaluations of weights and distances.
Working with 18 synthetic and real-life datasets, Wettschereck [50] found that reciprocally linear distance-weights clearly outperform unweighted NN for Euclidean distance, and that there is no clear difference between Euclidean and Boscovich distance. Zavrel [35], using cosine distance, additionally considered linear and Laplacian weights, but only linear weights clearly outperformed unweighted NN. Hechenbichler & Schliep [41] only evaluated linear and biquadratic weights, on a small number of datasets, without drawing any firm conclusions. However, we note that they appear to obtain generally better results for Boscovich than for Euclidean distance. Gou et al. [23], using Euclidean distance, found that their proposal outperforms both unweighted NN and linear distance-weights on twelve real-life datasets. Finally, the weights proposed by Gou et al. [23] also came out on top, along with linear and reciprocally quadratic distance-weights, in the extensive evaluations of time series classification conducted by Geler et al. [51, 52].
Table 1: Kernel functions $f$, with $a\in[0,1]$, $\lambda\in(-1,\infty)$, $p\in(0,\infty)$, and $m$ the number of features.
Name $f(a)$ Used in
Fuzzy negations
Linear $1-a$ [24, 25, 28, 30, 39, 41, 46]
Quadratic $1-a^{2}$ [30, 31]
Biquadratic $(1-a^{2})^{2}$ [41]
Samworth $1-a^{\frac{2}{m}}$ [22]
Sugeno-$\lambda$ $\frac{1-a}{1+\lambda\cdot a}$ [23]
Yager-$p$ $(1-a^{p})^{\frac{1}{p}}$
Other proper kernels
Constant $1$ [1, 39, 41, 45]
Laplace $e^{-a}$ [35, 39]
Gauss $e^{-a^{2}/2}$ [39]
Improper kernels
Reciprocally linear $\frac{1}{a}$ [28, 20, 32, 47]
Reciprocally quadratic $\frac{1}{a^{2}}$ [26, 20]
3 Proposals
In this section, we will discuss our novel proposals. These include a universal framework for nearest neighbour weighting for both NN and FRNN, a new weight type inspired by fuzzy Yager negation, an analysis of FNN, and a characterisation of scaling measures that relates them to distance measures.
3.1 Kernels
Figure 4: Proper kernel functions, expressed in terms of $a\in[0,1]$. In order to visualise the different weight that they place on smaller and larger values, we have rescaled each kernel by a constant, such that each kernel covers the same area on the interval $[0,1]$. As the resulting absolute values are essentially arbitrary, we have deliberately left the vertical axis unmarked.
In order to make nearest neighbour weighting easier to evaluate, we adopt a standardised representation based on kernel functions, inspired by the proposals discussed in Subsection 2.2.
The term kernel is used in various contexts for several related but distinct concepts. The type of kernel that we are interested in, from the statistical and signal processing literature, is a function on the real line that is symmetric around 0 as well as maximal around 0, such that it can be used to amplify values around 0 through multiplication. When such a kernel function has finite support, e.g. on $[-1,1]$, it can be used more specifically as a window function to filter out all values outside this range.
In the context of nearest neighbour weighting, all ranks and distance values are non-negative, and the weights are distributed over a finite range, limited by the $k$th rank or nearest neighbour distance. Therefore, all weight types that we will consider can be covered by the following more restricted definition of a kernel:
Definition 1.
A (proper) kernel is a decreasing function $f:[0,1]\longrightarrow\mathbb{R}_{\geq 0}$. A normalised kernel is a kernel with $f(0)=1$. An improper kernel is a decreasing function $f:(0,1]\longrightarrow\mathbb{R}_{\geq 0}$ with $\lim_{a\rightarrow 0}f(a)=\infty$. For a proper or improper kernel $f$, we write $f^{k}_{i}=f(\frac{i}{k+1})$.
Note that any proper kernel can be normalised through division by $f(0)$, such that $f(0)=1$ and the range of $f$ becomes $[0,1]$. An improper kernel is improper precisely because it cannot be normalised in this way.
A distinct concept from fuzzy set theory is that of a fuzzy negation (or fuzzy complement) [53], which generalises the ordinary negation from classical logic that sends 0 to 1 and 1 to 0:
Definition 2.
A fuzzy negation is a decreasing function $f:[0,1]\longrightarrow[0,1]$ with $f(0)=1$ and $f(1)=0$.
While these two concepts have different origins, we see that a fuzzy negation is in fact the same thing as a normalised kernel $f$ which satisfies $f(1)=0$. Conversely, we can view proper and improper kernels as a loose form of fuzzy negation. This correspondence allows us to interpret the effect of applying a kernel not just as placing greater weight on the contribution of nearby neighbours, but also as converting distance values into similarity values.
In the following subsections, we will redefine NN and FRNN classification in terms of kernel functions. The kernels used in this paper are listed in Table 1 and the proper kernels are visualised in Fig. 4.
3.2 NN
Using the definition of a kernel function allows us to state the following generalised definition for weighted nearest neighbour classification:
Definition 3.
Let $d$ be a distance measure, $k$ a positive integer, and $w$ and $s$ choices of kernel functions. Then the score for a decision class $C$ and a test record $y$ is:
$\left.\sum_{i\leq k|x_{i}\in C}w_{i}^{k}\cdot s(d_{i}^{*})\middle/\left(\sum_{i\leq k}w_{i}^{k}\cdot s(d_{i}^{*})\right)\right.,$
where $x_{i}$ is the $i$th nearest neighbour of $y$ in the training set $X$ (as determined by $d$), $d_{i}$ the corresponding distance and $d_{i}^{*}:=d_{i}/d_{k}$.
We adopt the following conventions to resolve specific edge cases:
  • 1.
    If $d_{k}=0$ (and therefore $d_{i}=0$ for all $i\leq k$), we stipulate $d_{i}^{*}:=d_{i}=0$.
  • 2.
    If $d_{1}=d_{k}$ (and therefore $d_{i}^{*}=1$ for all $i\leq k$) and if $s(1)=0$, we stipulate $s(d_{i}^{*}):=1$ for all $i$.
  • 3.
    If $s$ is an improper kernel, and $d_{i}^{*}=0$ for some $i$, we stipulate $s(d_{i}^{*}):=1$ for all such $i$ and $s(d_{i}^{*}):=0$ for all other $i$.
When $w$ is constant, we recover NN with distance-weights, when $s$ is constant, we recover NN with rank-weights, and when both $w$ and $s$ are constant, we recover unweighted NN. In addition, in all three edge cases listed above, we also effectively revert to performing unweighted classification with (part of) the nearest neighbours of $y$.
We will now show how this incorporates existing weighting proposals. Firstly, note that both in the original equation for NN (1), and in (3), we rescale each class score by the total sum of the weights. Therefore, we do not require that weights sum to 1. Moreover, multiplying all weights by a positive constant produces identical class scores. We will use the proportionality symbol $\propto$ to indicate that two sets of weights only differ by a positive constant factor. Thus, the linear distance-weights proposed by Dudani [28] can be simplified as follows:
$s_{i}$ $=\frac{d_{k}-d_{i}}{d_{k}-d_{1}}$
$\propto\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{d_{k}-d_{1}}{d_{k}}$
$=1-d_{i}^{*}.$
We see that Dudani’s linear weights are equivalent to applying a linear (or triangular) kernel in our revised definition. In fuzzy set theory, this is the original fuzzy negation introduced by both Łukasiewicz [54] and Zadeh [55].
In a similar way, we can simplify the weight types proposed by Gou et al. [23]:
$s_{i}$ $=\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{d_{k}+d_{1}}{d_{k}+d_{i}}$ (6)
$\propto\frac{d_{k}-d_{i}}{d_{k}-d_{1}}\cdot\frac{d_{k}+d_{1}}{d_{k}+d_{i}}\cdot\frac{d_{k}-d_{1}}{d_{k}+d_{1}}\cdot\frac{1}{d_{k}}$
$=\frac{1-d_{i}^{*}}{1+d_{i}^{*}}.$
We will call this the Sugeno-1 kernel, because it is in fact an instance of fuzzy Sugeno negation [56], with $\lambda=1$:
$a\longmapsto\frac{1-a}{1+\lambda a}$
Weights that depend reciprocally linearly or reciprocally quadratically on nearest neighbour distance result from a straightforward application of an improper kernel, since, respectively,
$s_{i}=\frac{1}{d_{i}}\propto\frac{d_{k}}{d_{i}}=\frac{1}{d_{i}^{*}},$
and
$s_{i}=\frac{1}{d_{i}^{2}}\propto\frac{d_{k}^{2}}{d_{i}^{2}}=\frac{1}{\left(d_{i}^{*}\right)^{2}}.$
We can also simplify the weight types proposed by Macleod et al. [37]:
$s_{i}$ $=\frac{d_{k}-d_{i}+d_{k}-d_{1}}{2(d_{k}-d_{1})}$
$\propto\frac{d_{k}-d_{i}+d_{k}-d_{1}}{2(d_{k}-d_{1})}\cdot\frac{2(d_{k}-d_{1})}{d_{k}}$
$=2-d_{i}^{*}-d_{1}^{*}.$
However, the resulting function still depends on $d_{1}$. That means that we can use this function to calculate distance-weights, but it does not generalise to a kernel that we could also apply to rank-weights.
The only remaining distance-weights that cannot be rewritten into a kernel function are the Laplace weights $e^{-d_{i}}$ proposed in [35]. Note that these are not homogeneous, i.e. the weighting depends on the absolute scale of the distances, which is arguably undesirable. However, we can consider the Laplace kernel $e^{-d_{i}^{*}}$.
Similarly, most types of rank-weights proposed in the literature can be obtained by applying a kernel function to $\frac{i}{k+1}$. For the Samworth weights, if we fix a particular value $k>0$, and write $h=\frac{1}{k+1}$ (for reasons of space), we can define the following kernel function $f^{k}$ such that $f^{k}(\frac{i}{k+1})$ is the $i$th weight as in (2):
$f^{k}(a)=\frac{1}{k}\left(1+\frac{m}{2}-\frac{m}{2k^{\frac{2}{m}}}\left((a/h)^{1+\frac{2}{m}}-(a/h-1)^{1+\frac{2}{m}}\right)\right).$
$f^{k}$ is proportional to $\frac{2k}{m+2}\cdot f^{k}$, and we have the following lemma:
Lemma 1.
$\lim_{k\rightarrow\infty}\frac{2k}{m+2}\cdot f^{k}=1-a^{\frac{2}{m}}.$
Proof.
$\mathrel{\phantom{=}}\lim_{k\rightarrow\infty}\frac{2k}{m+2}\cdot f_{k}$
$=\lim_{k\rightarrow\infty}1-\frac{m}{m+2}\cdot\frac{1}{k^{\frac{2}{m}}}\left((a/h)^{1+\frac{2}{m}}-(a/h-1)^{1+\frac{2}{m}}\right)$
$=1-\frac{m}{m+2}\cdot\lim_{k\rightarrow\infty}\frac{(a/h)^{1+\frac{2}{m}}-(a/h-1)^{1+\frac{2}{m}}}{k^{\frac{2}{m}}}$
$=1-\frac{m}{m+2}\cdot\lim_{k\rightarrow\infty}\frac{h^{1+\frac{2}{m}}}{h^{1+\frac{2}{m}}}\frac{(a/h)^{1+\frac{2}{m}}-(a/h-1)^{1+\frac{2}{m}}}{k^{\frac{2}{m}}}$
$=1-\frac{m}{m+2}\cdot\lim_{k\rightarrow\infty}\frac{a^{1+\frac{2}{m}}-(a-h)^{1+\frac{2}{m}}}{h\cdot\frac{k^{2/m}}{(k+1)^{2/m}}}$
$=1-\frac{m}{m+2}\cdot\lim_{h\rightarrow 0}\frac{a^{1+\frac{2}{m}}-(a-h)^{1+\frac{2}{m}}}{h}$
$=1-\frac{m}{m+2}\cdot\left(1+\frac{2}{m}\right)a^{\frac{2}{m}}$ (*)
$=1-a^{\frac{2}{m}},$
where ()(*)( ∗ ) is the polynomial rule for derivation. ∎
Accordingly, we call $1-a^{\frac{2}{m}}$ the Samworth kernel.
The only rank-weights that can not be obtained by applying a kernel function are the Fibonacci weights from [33], because their relative distribution depends on $k$.
By reformulating both distance-weights and rank-weights in terms of a kernel function, we obtain a single unique way to characterise all the different weight types. In addition, this representation makes it clear that we could also choose to apply e.g. the Sugeno-1 kernel to obtain rank-weights, or the Samworth kernel to obtain distance-weights, even though they were originally proposed for, respectively, distance-weights and rank-weights. Furthermore, we could choose to apply both rank- and distance-weights at the same time, as in the proposal by Gou et al. [32], which can be realised by combining a reciprocally linear rank-kernel and a linear distance-kernel.
3.3 Yager weights
In the previous section we found that two weighting proposals from the literature correspond, respectively, to classical and Sugeno negation. There is a third type of fuzzy negation that is frequently encountered in the literature, with the following form:
$a\longmapsto(1-a^{p})^{\frac{1}{p}},$
for some $p>0$. This is generally known as Yager negation, because it was proposed by Higashi & Klir [53] to accompany other operators introduced by Yager [57].
Just like classical and Sugeno negation, Yager negation can be used for nearest neighbour weighting. Specifically, we propose to use Yager negation with $p=\frac{1}{2}$, because the resulting contour is quite different from that of the existing weighting proposals, except the Samworth kernel for larger values of $m$ (Fig. 4).
3.4 FNN
Recall that in the original proposal of FNN (Subsection 2.3), there were two possible values for $u_{i}$. When $u_{i}$ is chosen crisply, (4) simplifies to
$\left.\sum\limits_{i\leq k|x_{i}\in C}1/d_{i}^{2/(q-1)}\middle/\sum\limits_{i\leq k}1/d_{i}^{2/(q-1)}\right..$
In other words, FNN becomes equivalent to NN classification (1), with
$s_{i}=1/d_{j}^{\frac{2}{q-1}},$
for some $q>1$. When $q=3$ and $q=2$, we obtain, respectively, reciprocally linear and reciprocally quadratic distance-weights.
Therefore, we will continue with the fuzzy variant of $u_{i}$ (5). In that case, (4) becomes:
$\mathrel{\phantom{=}}\left.\sum\limits_{i\leq k}u_{i}\cdot s_{i}\middle/\sum\limits_{i\leq k}s_{i}\right.$
=(${\sum_{i\leq k|x_{i}\in C}^{}}0.51\cdot s_{i}+\sum\limits_{i\leq k}0.49\cdot n_{C}(x_{i})/k\cdot s_{i}\right)\middle/\sum_{i\leq k}s_{i}\right.$
$=\left.0.51\cdot\sum_{\mathclap{i\leq k|x_{i}\in C}}s_{i}\middle/\sum_{i\leq k}s_{i}\right.+\left.0.49\cdot\sum_{i\leq k}n_{C}(x_{i})/k\cdot s_{i}\middle/\sum_{i\leq k}s_{i}\right.$
Thus, in this variant, the FNN class score is the weighted average of two components. The first component is, again, NN, while the second component is NN with fuzzified class membership. This second component can be rewritten as:
$\mathrel{\phantom{=}}\left.\sum_{i\leq k}n_{C}(x_{i})/k\cdot s_{i}\middle/\sum_{i\leq k}s_{i}\right.$
$=\left.\frac{1}{k}\sum_{i,j\leq k|x_{ij}\in C}s_{i}\middle/\sum_{i\leq k}s_{i}\right.$
where $x_{ij}$ is the $j$th neighbour of the $i$th neighbour of $y$. In effect, this is also NN, with class scores that are on the one hand diluted (being based not just on the class scores of the $k$ nearest neighbours of $y$, but also on the class scores of their $k$ nearest neighbours) and on the other hand concentrated (because nearby neighbours will more frequently appear as neighbours of neighbours). The net effect of this on classification performance remains to be evaluated experimentally, but it is hard to see how this reduces the influence of ‘atypical’ training records, the original stated motivation for FNN.
Table 2: Real-life Classification Datasets from the UCI Repository for Machine Learning. $n$: number of records; $c$: number of classes; $m$: number of attributes; IR: imbalance ratio.
Dataset $n$ $c$ $m$ IR Dataset $n$ $c$ $m$ IR
accent 329 6 12 2.5 mfeat 2000 10 649 1.0
acoustic-features 400 4 50 1.0 miniboone 130 064 2 50 2.6
ai4i2020 10 000 2 6 28.5 new-thyroid 215 3 5 3.5
alcohol 125 5 12 1.0 oral-toxicity 8992 2 1024 11.1
androgen-receptor 1687 2 1024 7.5 page-blocks 5473 5 10 31.6
avila 20 867 12 10 38.7 phishing-websites 11 055 2 30 1.3
banknote 1372 2 4 1.2 plrx 182 2 12 2.5
bioaccumulation 779 3 9 4.3 pop-failures 540 2 18 10.7
biodeg 1055 2 41 2.0 post-operative 87 2 8 2.6
breasttissue 106 6 9 1.3 qualitative-bankruptcy 250 2 6 1.3
ca-cervix 72 2 19 2.4 raisin 900 2 7 1.0
caesarian 80 2 5 1.4 rejafada 1996 2 6824 1.0
ceramic 37 4 34 1.4 rice 3810 2 7 1.3
cmc 1473 3 9 1.6 seeds 210 3 7 1.0
codon-usage 13 011 20 64 25.5 segment 2310 7 19 1.0
coimbra 116 2 9 1.2 seismic-bumps 2584 2 18 14.2
column 310 3 6 1.9 sensorless 58 509 11 48 1.0
debrecen 1151 2 19 1.1 sepsis-survival 110 204 2 3 12.6
dermatology 358 6 34 2.2 shuttle 58 000 7 9 560.8
diabetes-risk 520 2 16 1.6 skin 245 057 2 3 3.8
divorce 170 2 54 1.0 somerville 143 2 6 1.2
dry-bean 13 611 7 16 2.3 sonar 208 2 60 1.1
ecoli 332 6 7 6.3 south-german-credit 1000 2 20 2.3
electrical-grid 10 000 2 12 1.8 spambase 4601 2 57 1.5
faults 1941 7 27 3.9 spectf 267 2 44 3.9
fertility 100 2 9 7.3 sportsarticles 1000 2 59 1.7
flowmeters 361 4 44 1.7 sta-dyn-lab 6248 2 244 9.5
forest-types 523 4 9 1.8 tcga-pancan-hiseq 801 5 20 531 1.9
gender-gap 3145 2 15 7.9 thoraric-surgery 470 2 16 5.7
glass 214 6 9 3.6 transfusion 748 2 4 3.2
haberman 306 2 3 2.8 tuandromd 4464 2 241 4.0
hcv 589 2 12 9.5 urban-land-cover 675 9 147 2.2
heart-failure 299 2 12 2.1 vehicle 846 4 18 1.1
house-votes-84 435 2 16 1.6 warts 180 2 8 2.0
htru2 17 898 2 8 9.9 waveform 5000 3 21 1.0
ilpd 579 2 10 2.5 wdbc 569 2 30 1.7
ionosphere 351 2 34 1.8 wifi 2000 4 7 1.0
iris 150 3 4 1.0 wilt 4839 2 5 17.5
landsat 6435 6 36 1.7 wine 178 3 13 1.3
leaf 340 30 14 1.2 wisconsin 683 2 9 1.9
letter 20 000 26 16 1.0 wpbc 138 2 32 3.9
lrs 527 7 100 12.6 yeast 1484 10 8 11.6
magic 19 020 2 10 1.8
3.5 FRNN
The upper and lower approximations of FRNN can similarly be rewritten using kernel functions:
Definition 4.
Let $d$ be a distance measure and $k$ a positive integer, $w$ a choice of kernel function and $s$ a choice of fuzzy negation. Then the score for a decision class $C$ and a test record $y$ is:
[…]
where $d^{+}_{i}$ and $d^{-}_{i}$ are the $i$th nearest neighbour distance of $y$ in, respectively, $C$ and $X\setminus C$, and $d^{+}_{*}$ and $d^{-}_{*}$ are to be defined.
$d_{*}^{+}$ and $d_{*}^{-}$ determine cutoff values — all larger distances are mapped to the minimum degree of similarity, typically 0. These cutoff values have to be constant across decision classes and test records, to allow for a proper comparison of class scores. If we choose values that are too small, $\min(d_{i}^{+}/d_{*}^{+},1)$ and $\min(d_{i}^{-}/d_{*}^{-},1)$ become equal to 1 for many test records and many values of $i$, and we lose information. If we choose values that are too large, $\min(d_{i}^{+}/d_{*}^{+},1)$ and $\min(d_{i}^{-}/d_{*}^{-},1)$ are generally close to 00, and we do not make full use of the profile of the kernel $s$. Therefore, as a compromise, we calculate $d_{k}^{+}$ and $d_{k}^{-}$ of all training records, for all decision classes, and take $d_{*}^{+}$ and $d_{*}^{-}$ to be the respective maxima of these values.
Note that unlike NN, FRNN cannot be used with constant distance-weights, because this would equalise all class scores. Instead, the default choice is linear distance-weights, in which case the double negation $(1-s(d_{i}))$ in the lower approximation simplifies to $d_{i}$. In addition, the exponentially decreasing rank-weights that have occasionally been proposed in the literature have a very limited usefulness, as the contribution of each additional value quickly becomes insignificant, and, eventually, impossible to compute.
3.6 Distance and scaling measures
Three distance measures that are frequently used with nearest neighbour classification are Euclidean, Boscovich (or city-block) and Chebyshev (or maximum) distance. These can be viewed, respectively, as the special cases $p=2$, $p=1$ and $p\rightarrow\infty$ of the Minkowski $p$-distance between two points $x,y\in\mathbb{R}^{m}$ (for some $m\geq 1$):
$\left|y-x\right|_{p}:=\left(\sum_{i\leq m}\left|y_{i}-x_{i}\right|^{p}\right)^{\frac{1}{p}}.$
In order to obtain a comparable contribution from all attributes, these must be rescaled to a common scale. This can be done by taking a measure of dispersion, and dividing each attribute by this measure, such that it becomes 1 for each attribute. Two common choices are the standard deviation and the half-range of each attribute. These can be linked to the concept of Minkowski $p$-distance by defining the Minkowski $p$-centre and $p$-radius of a dataset:
Definition 5.
Let $X=(x_{1},x_{2},\dots,x_{n})$ be a univariate real-valued dataset. The Minkowski $p$-radius $r_{p}$ of $X$ is defined as:
$r_{p}(X):=\min_{z\in\mathbb{R}}\left(\frac{1}{n}\sum_{i\leq n}\left|x_{i}-z\right|^{p}\right)^{\frac{1}{p}},$
while the Minkowski $p$-centre of $X$ is the corresponding minimising value for $z$ (not necessarily unique for $p\leq 1$).
The standard deviation and half-range of a dataset are $r_{2}$ and $r_{\infty}$, while the corresponding $2$-centre and $\infty$-centre of a dataset are its mean and its midrange. The $1$-centre of a dataset is its median, and the corresponding measure of dispersion $r_{1}$ that it minimises is the mean absolute deviation around the median. Thus, $r_{1}$ is another measure of dispersion that we can use to scale attributes with.
A potential advantage of $r_{1}$-scaling over $r_{2}$-scaling is its reduced sensitivity to outliers, as $r_{1}$ only depends linearly on outliers, rather than quadratically like $r_{2}$. In turn, both measures are much less sensitive to outliers than $r_{\infty}$, which is completely determined by the most extreme outlier. An alternative way to obtain a measure of dispersion that is less sensitive to outliers is to explicitly ignore peripheral values. Specifically, we will consider the semi-interquartile range $r_{\infty}^{*}$, which is the half-range of the central 50% of all values.
4 Experimental setup
To evaluate NN, FNN and FRNN classification, we will use 85 numerical real-life datasets from the UCI repository for machine learning (Table 2). We perform 5-fold cross-validation, and calculate the mean area under the receiver operating characteristic (AUROC) as a measure of the discriminative ability of each classifier. For FRNN, this requires normalising class scores such that they sum to 1. To compare two alternatives, we calculate the $p$-value from a one-sided Wilcoxon signed-rank test. Where appropriate, we also apply the Holm-Bonferroni method [58] to correct for family-wise error.
For all of NN, FNN and FRNN, we optimise $k$ through leave-one-out validation, evaluating all values up to a certain value $k^{\max}$. For FRNN, we also choose between the upper, lower or mean approximation based on validation AUROC.
Determining the best validation procedure for hyperparameter selection is a surprisingly complex issue. As estimators of the test set score, validation procedures suffer from both bias and variance. Compared to e.g. 10-fold cross-validation, leave-one-out validation has lower bias, but can have higher variance when the model is unstable [59], which depends on the classification procedure, the evaluation measure and the structure of the dataset, in particular the amount of outliers [60]. Complicating things further, a validation procedure that is worse for estimating prediction performance can still, somewhat counter-intuitively, be better for selecting the optimal hyperparameter values, and vice-versa [59].
In our case, there are three particular reasons that justify the use of leave-one-out validation. Firstly, nearest neighbour classification is very stable [61]. Secondly, the optimal value of the hyperparameter that we optimise ($k$) is likely influenced by the dataset size, so it is advantageous that the internal training set size of leave-one-out validation ($n-1$) is nearly the same as that of the full training set ($n$). And thirdly, leave-one-out validation can be performed very efficiently for nearest neighbour algorithms, by executing a single $(k+1)$-nearest neighbour query on the training set and deleting every match between a training record and itself.
In addition to the use of leave-one-out validation, there is a second reason why $k$ can be optimised very efficiently: we do not need to perform a new nearest neighbour query for every value of $k$ that we want to evaluate. Instead, we only need to perform a single $k^{\max}$-nearest neighbour query, after which we can evaluate all values of $k$ from $k^{\max}$ to 1 by iteratively deleting all $(k+1)$-nearest neighbour distances.
Because larger datasets may require larger values of $k$, we let $k^{\max}$ depend logarithmically on the training set size $n$. In order to ensure high model quality, we set $k^{\max}=100\log n$. For our largest dataset, skin, this means that $k^{\max}=1219$ in each fold.
5 Results
In this section, we will present the results of our experiments. To start with, we will evaluate distance measures, scaling measures and weight types, but restrict ourselves to the weight types that have previously been proposed in the literature. We will then ask whether these results can be further improved upon by using the Yager-$\frac{1}{2}$ weights that we have proposed.
5.1 NN
Table 3: One-sided $p$-values, Samworth distance-weights vs other distance-weights, for NN with Boscovich distance, in terms of AUROC. Holm-Bonferroni family-wise error correction applied to each column.
Samworth vs… Scaling
$r_{1}$ $r_{2}$ $r_{\infty}$ $r_{\infty}^{*}$
Constant $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Quadratic $0.00051$ $0.00016$ $<0.0001$ $<0.0001$
Gauss $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Laplace $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Linear $0.00016$ $0.00030$ $0.00048$ $<0.0001$
Macleod $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Biquadratic $0.014$ $0.11$ $0.71$ $0.042$
Reciprocally linear $0.0012$ $0.00016$ $<0.0001$ $<0.0001$
Reciprocally quadratic $0.038$ $0.020$ $0.046$ $0.059$
Sugeno-1 $0.0068$ $0.0036$ $0.029$ $0.00036$
We first consider the effect of the distance on classification performance. We find that for all types of scaling and all weight types, Boscovich distance leads to significantly better performance than Euclidean ($p<0.0031$) and Chebyshev ($p<2.0\cdot 10^{-9}$) distance. For this reason, we will only consider Boscovich distance for the rest of our analysis.
Next, we have a look at the different kernels. First we compare using each kernel for distance-weights versus rank-weights, for each type of scaling. With three exceptions, we find that distance-weights lead to significantly better classification performance than rank-weights ($p<0.0019$). The exceptions are the reciprocal and Laplace kernels, for which the difference is not or only weakly significant and to which we return below, as well as the Gaussian kernel, for which the difference is only weakly significant with $r_{\infty}$-scaling ($p=0.26$).
Among the distance-weights, the Samworth kernel significantly outperforms all other kernels (Table 3), with three exceptions. The difference with respect to the biquadratic and reciprocally quadratic kernel is only weakly significant with, respectively, $r_{2}$- and $r_{\infty}^{*}$-scaling. Moreover, with $r_{\infty}$-scaling, the biquadratic kernel is actually slightly better than the Samworth kernel on our data, but we will see below that $r_{\infty}$-scaling is suboptimal. Finally, we also find that Samworth distance-weights outperform reciprocal ($p<0.0048$) and Laplace ($p<6.2\cdot 10^{-6}$) rank-weights across scaling measures.
Samworth distance-weights also significantly outperform the combination of linear distance-weights and reciprocal rank-weights proposed by Gou et al. [32] for all scaling types ($p<0.044$) except $r_{\infty}$, where the difference is only weakly significant ($p=0.13$). However, our general formula for NN classification also allows for other combinations. Indeed, we find that Samworth distance-weights are outperformed by the combination of Samworth distance-weights and Samworth rank-weights ($p<0.040$).
Table 4: One-sided $p$-values, various scalings vs $r_{\infty}$-scaling, for NN with Boscovich distance and Samworth distance- and rank-weights, in terms of AUROC. Holm-Bonferroni family-wise error correction applied.
Test p
$r_{2}$ vs $r_{\infty}$ $0.00044$
$r_{1}$ vs $r_{\infty}$ $0.023$
$r_{\infty}^{*}$ vs $r_{\infty}$ $0.023$
Finally, when we consider the different measures of dispersion that can be used to normalise a dataset through rescaling, we find that $r_{1}$, $r_{2}$ and $r_{\infty}^{*}$ do not significantly outperform each other for the combination of Samworth distance- and rank-weights, while they all outperform $r_{\infty}$ (Table 4). For other weight types, we obtain comparable results.
5.2 FNN
For FNN, we consider reciprocally linear and reciprocally quadratic distance-weights, as well as Samworth distance-weights and a combination of Samworth rank- and distance-weights, since we found in the previous Subsection that these latter two perform well for classical NN.
As with NN, FNN performs significantly better with Boscovich distance than with either Euclidean ($p<0.025$) or Chebyshev ($p<8.3\cdot 10^{-7}$) distance.
Unlike NN, it is not clear that Samworth distance-weights perform better than reciprocally linear or reciprocally quadratic weights (Table 5). Furthermore, the combination of Samworth rank- and distance-weights actually performs worse than Samworth distance-weights alone for $r_{1}$ ($p=0.032$) and $r_{\infty}$ ($p=0.000078$) scaling, while for $r_{2}$- and $r_{\infty}^{*}$-scaling, the difference is not significant. We have weak evidence that with FNN, $r_{1}$-scaling leads to better performance than $r_{2}$-scaling, and that in turn both are preferable over $r_{\infty}$- and $r_{\infty}^{*}$-scaling (Table 6).
Table 5: One-sided $p$-values, Samworth distance-weights vs other distance-weights, for FNN with Boscovich distance, in terms of AUROC. Holm-Bonferroni family-wise error correction applied to each column.
Samworth vs… Scaling
$r_{1}$ $r_{2}$ $r_{\infty}$ $r_{\infty}^{*}$
Reciprocally linear $0.028$ $0.17$ $0.012$ $0.68$
Reciprocally quadratic $0.34$ $0.32$ $0.18$ $0.69$
Table 6: One-sided $p$-values, comparing the scaler in each row against the scaler in each column, for FNN with Boscovich distance and Samworth distance-weights, in terms of AUROC. Holm-Bonferroni family-wise error correction applied to each row.
$r_{2}$ $r_{\infty}$ $r_{\infty}^{*}$
$r_{1}$ $0.19$ $0.14$ $0.14$
$r_{2}$ $0.21$ $0.21$
$r_{\infty}$ $0.98$
Table 7: One-sided $p$-values, Samworth vs linear distance-weights, for FRNN with Boscovich distance and various rank-weights and scalings.
Rank-kernel Scaling
$r_{1}$ $r_{2}$ $r_{\infty}$ $r_{\infty}^{*}$
Constant $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Linear $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Reciprocally linear $0.082$ $0.12$ $0.061$ $0.0061$
Samworth $0.0012$ $0.00019$ $0.00043$ $<0.0001$
Table 8: One-sided $p$-values, Samworth rank-weights vs other rank-weights, for FRNN with Boscovich distance and Samworth distance-weights. Holm-Bonferioni correction applied in each column.
Samworth vs… Scaling
$r_{1}$ $r_{2}$ $r_{\infty}$ $r_{\infty}^{*}$
Constant $0.18$ $0.0033$ $0.089$ $0.0048$
Linear $0.18$ $0.074$ $0.73$ $0.32$
Reciprocally linear $0.18$ $0.022$ $0.15$ $0.057$
5.3 FRNN
Table 9: One-sided $p$-values, various scalings vs $r_{\infty}$-scaling, for FRNN with Boscovich distance and Samworth distance- and rank-weights, in terms of AUROC. Holm-Bonferroni family-wise error correction applied.
Test p
$r_{2}$ vs $r_{\infty}$ $0.0078$
$r_{1}$ vs $r_{\infty}$ $0.060$
$r_{\infty}^{*}$ vs $r_{\infty}$ $0.060$
For FRNN, we evaluate the different types of rank-weights proposed in the literature, corresponding to the constant, linear and reciprocal kernel, in combination with linear distance-weights. In addition, we evaluate Samworth rank- and distance-weights, motivated by their excellent performance with NN.
Here we also find that Boscovich distance leads to higher AUROC than Euclidean ($p<0.00085$) and Chebyshev ($p<6.2\cdot 10^{-10}$) distance for all combinations of kernels and scaling measures.
The traditional choice for distance-weights is to use a linear kernel, but we find that the Samworth kernel performs better, although the difference is only weakly significant in combination with reciprocal rank-weights (Table 7).
Likewise, Samworth rank-weights appear to be the best choice, but the advantage over other kernels is only weakly significant (Table 8).
As with NN, the measures of dispersion $r_{1}$, $r_{2}$ and $r_{\infty}^{*}$ do not significantly outperform each other, but do outperform $r_{\infty}$, although even this latter fact is only weakly significant (Table 9).
5.4 NN vs FNN vs FRNN
Table 10: One-sided $p$-values, NN vs FNN with Boscovich distance and various distance kernels and scalings, in terms of AUROC.
Distance-kernel Scaling
$r_{1}$ $r_{2}$ $r_{\infty}$ $r_{\infty}^{*}$
Reciprocally linear $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Reciprocally quadratic $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Samworth $<0.0001$ $<0.0001$ $<0.0001$ $<0.0001$
Table 11: One-sided $p$-values, FRNN vs NN, Boscovich distance and Samworth rank- and distance-weights.
Scaling
$r_{1}$ $r_{2}$ $r_{\infty}$ $r_{\infty}^{*}$
$0.021$ $0.0092$ $0.0052$ $0.049$
Table 12: Mean 5-fold cross-validation AUROC with Boscovich distance, $r_{2}$-scaling and Samworth distance-weights (FNN) or both Samworth distance- and rank-weights (NN, FRNN). Bold: highest value (before rounding).
dataset NN FNN FRNN dataset NN FNN FRNN
accent 0.976 0.973 0.971 mfeat 0.999 0.999 1.000
acoustic-features 0.930 0.921 0.938 miniboone 0.962 0.954 0.962
ai4i2020 0.957 0.911 0.954 new-thyroid 0.989 0.985 0.994
alcohol 1.000 1.000 1.000 oral-toxicity 0.888 0.871 0.878
androgen-receptor 0.855 0.847 0.867 page-blocks 0.975 0.966 0.982
avila 0.986 0.979 0.987 phishing-websites 0.996 0.994 0.994
banknote 1.000 1.000 1.000 plrx 0.479 0.442 0.496
bioaccumulation 0.741 0.733 0.755 pop-failures 0.947 0.949 0.941
biodeg 0.926 0.903 0.926 post-operative 0.476 0.478 0.479
breasttissue 0.932 0.902 0.921 qualitative-bankruptcy 1.000 0.995 1.000
ca-cervix 0.949 0.910 0.949 raisin 0.927 0.909 0.925
caesarian 0.708 0.708 0.685 rejafada 0.958 0.950 0.957
ceramic 0.758 0.765 0.800 rice 0.979 0.964 0.979
cmc 0.701 0.627 0.701 seeds 0.991 0.979 0.991
codon-usage 0.969 0.967 0.975 segment 0.998 0.997 0.998
coimbra 0.774 0.760 0.794 seismic-bumps 0.767 0.673 0.773
column 0.919 0.905 0.915 sensorless 1.000 0.999 1.000
debrecen 0.734 0.676 0.734 sepsis-survival 0.700 0.701 0.666
dermatology 0.999 0.998 0.999 shuttle 0.998 0.997 1.000
diabetes-risk 0.998 0.991 0.999 skin 1.000 1.000 1.000
divorce 1.000 1.000 1.000 somerville 0.595 0.561 0.614
dry-bean 0.996 0.994 0.994 sonar 0.959 0.927 0.961
ecoli 0.969 0.947 0.972 south-german-credit 0.789 0.756 0.794
electrical-grid 0.968 0.959 0.968 spambase 0.980 0.977 0.977
faults 0.959 0.950 0.962 spectf 0.841 0.837 0.852
fertility 0.687 0.763 0.647 sportsarticles 0.880 0.861 0.884
flowmeters 0.977 0.961 0.970 sta-dyn-lab 0.994 0.992 0.996
forest-types 0.969 0.958 0.974 tcga-pancan-hiseq 1.000 1.000 1.000
gender-gap 0.704 0.623 0.700 thoraric-surgery 0.624 0.545 0.630
glass 0.942 0.917 0.948 transfusion 0.725 0.689 0.747
haberman 0.681 0.633 0.709 tuandromd 0.997 0.998 0.999
hcv 0.993 0.992 0.994 urban-land-cover 0.973 0.968 0.974
heart-failure 0.867 0.863 0.863 vehicle 0.912 0.887 0.913
house-votes-84 0.984 0.979 0.983 warts 0.907 0.870 0.899
htru2 0.972 0.963 0.977 waveform 0.972 0.971 0.971
ilpd 0.727 0.675 0.734 wdbc 0.989 0.989 0.994
ionosphere 0.942 0.935 0.982 wifi 1.000 0.999 0.999
iris 0.997 0.998 0.997 wilt 0.962 0.940 0.974
landsat 0.990 0.987 0.989 wine 1.000 1.000 1.000
leaf 0.975 0.961 0.975 wisconsin 0.990 0.991 0.996
letter 0.999 0.999 0.999 wpbc 0.589 0.573 0.665
lrs 0.915 0.885 0.899 yeast 0.869 0.860 0.890
magic 0.916 0.893 0.923
In Subsection 5.2, we observed that FNN performs best on our data with Samworth distance-weights, but that the difference with respect to reciprocally linear and reciprocally square distance-weights is not significant for all scaling-types. However, when we compare FNN to NN, we find that for all three kernels, NN performs significantly better (Table 10).
For both NN and FRNN, we obtained the best results with a combination of Samworth distance- and rank-weights. When we compare NN and FRNN against each other, we find that FRNN performs better (Table 11).
Table 12 lists the mean AUROC achieved with NN, FNN and FRNN for each dataset with $r_{2}$-scaling (the results are comparable for the other scaling types). Fig.5 illustrates the increase in performance for NN and FRNN due to the various hyperparameter choices proposed in this paper.
Figure 5: Violin plots of the distribution of AUROC scores across our selection of datasets, illustrating the cumulative benefit of choosing optimal hyperparameter values. We start with A: choosing $k=1$, Euclidean distance and $r_{\infty}$-scaling, and successively change this by B: choosing $k=5$ with constant weights (unweighted); C: setting $k$ and the approximation type (for FRNN) through leave-one-out validation; D: choosing Boscovich distance; E: choosing Samworth distance- and rank-weights; and F: choosing $r_{2}$-scaling. Dashed lines: median; dotted lines: first and third quartiles.
5.5 Yager-$\frac{1}{2}$ weights
Finally, we consider the results of the Yager-$\frac{1}{2}$ kernel that we have proposed. When we equip NN with both Yager-$\frac{1}{2}$ distance- and rank-weights, this performs slightly better on our data than Samworth distance- and rank-weights, but the difference is not significant ($p<0.25$ across scaling measures). Interestingly, unlike the Samworth kernel, the Yager-$\frac{1}{2}$ kernel appears to perform about as well when only used for distance-weights and when used for both distance- and rank-weights. Correspondingly, Yager-$\frac{1}{2}$ distance-weights perform significantly better than Samworth distance-weights ($p<0.0053$). Thus the main advantage of the Yager-$\frac{1}{2}$ kernel is that it enables comparable performance as the Samworth kernel, but is easier to implement, because it does not require the addition of rank-weights and because it is not dependent on the dimensionality of the dataset.
In contrast, for FNN we obtain comparable performance between Samworth and Yager-$\frac{1}{2}$ distance-weights with $r_{1}$- or $r_{2}$-scaling, and for FRNN we find that Samworth distance- and rank-weights still perform significantly better than Yager-$\frac{1}{2}$ distance- and rank-weights ($p<0.050$ across scaling measures).
6 Conclusion
In this paper, we have provided a comprehensive overview of the different weighting variants of NN, FNN and FRNN classification that have been proposed in the literature. We have proposed a uniform framework for these proposals and conducted an evaluation on 85 real-life datasets. This allows us to draw the following conclusions:
  • 1.
    Weighting can be expressed as the application of a kernel function to the distances and/or ranks of the nearest neighbours of a test record — we have provided an overview of kernel functions that correspond to existing weighting proposals in the literature.
  • 2.
    In particular, Samworth rank-weights, which have been shown to be theoretically optimal, converge to a kernel function that depends on the dimensionality of the data, and that can also be applied to obtain distance-weights.
  • 3.
    On real-life datasets, both NN and FRNN perform better with a combination of Samworth rank- and distance-weights than with other weight types proposed in the literature, while FNN appears to perform best with Samworth distance-weights and constant rank-weights.
  • 4.
    However, NN and FNN appear to perform equally well with Yager-$\frac{1}{2}$ weights, a novel weight type inspired by fuzzy Yager negation. For NN, the Yager-$\frac{1}{2}$ kernel offers two practical benefits over the Samworth kernel: it only needs to be applied to obtain distance-weights (not rank-weights), and it does not depend on the dimensionality of the dataset. For FRNN, Samworth weights still perform better.
  • 5.
    Boscovich distance clearly outperforms Euclidean and Chebyshev distance, regardless of other hyperparameter choices.
  • 6.
    With Samworth and Yager-$\frac{1}{2}$ weights, rescaling attributes by $r_{1}$ (mean absolute deviation around the median), $r_{2}$ (standard deviation) or $r_{\infty}^{*}$ (semi-interquartile range) produces comparable results, while these are all better than rescaling by $r_{\infty}$ (half-range).
  • 7.
    Our comparison between NN and FNN with identical distance-weights reveals that in practice, the fuzzification of class membership degrees in FNN leads to systematically lower performance. In contrast, with its more fundamentally different approach, FRNN does generally outperform NN when both are equipped with their best-performing weighting scheme (Samworth distance- and rank-weights).
We believe that these results serve as a useful baseline for future applications and research. For applications, we recommend the use of FRNN classification with Samworth rank- and distance-weights, Boscovich distance, and any one of $r_{1}$-, $r_{2}$- or $r_{\infty}^{*}$-scaling, while $k$ can be optimised through efficient leave-one-out validation. With the classical NN algorithm, we recommend the same hyperparameter choices, except that Yager-$\frac{1}{2}$ distance-weights may be substituted and rank-weights omitted.
We suggest that future research should concentrate on identifying even better-performing kernel functions. For this, the contours of the Samworth and Yager-$\frac{1}{2}$ kernels may serve as a useful starting point. In particular, we hope that in this way the following two questions may be answered:
  • 1.
    Should an optimal kernel depend on the dimensionality of the data? The good performance of the Samworth kernel suggests that the answer is yes. However, we note that its profile is actually very similar for typical dimensionalities of $4\leq m\leq 256$, so that its variation according to dimensionality may be far less important than the basic outline of its profile. This is seemingly confirmed by the relatively strong performance of the Yager-$\frac{1}{2}$ kernel, which has a similar profile.
  • 2.
    For NN, do we need both distance- and rank-weights? The fact that Yager-$\frac{1}{2}$ distance-weights alone perform as well as a combination of Samworth distance- and rank-weights suggests that with the right kernel, we can do away with rank-weights.
We hope that any future evaluations of competing proposals will be based on a similar collection of real-life classification problems to the one that we have assembled, which we have made available for this purpose.
CRediT authorship contribution statement
Oliver Urs Lenz: Conceptualisation, Software, Formal analysis, Investigation, Data Curation, Writing — Original Draft. Henri Bollaert: Conceptualisation, Investigation, Writing — Review & editing. Chris Cornelis: Writing — Review & editing, Supervision.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The datasets used in this paper can be downloaded from https://cwi.ugent.be/~oulenz/datasets/lenz-2024-unified.tar.gz.
Acknowledgments
The research reported in this paper was conducted with the financial support of the Odysseus programme of the Research Foundation – Flanders (FWO).
This publication is part of the project Digital Twin with project number P18-03 of the research programme TTW Perspective, which is (partly) financed by the Dutch Research Council (NWO).
We thank Martine De Cock for raising the important question whether NN and FRNN weighting could be unified, which sparked our inspiration for the present framework.